OpenStudy (anonymous):

How many atoms of 12C are in a 3.50 g sample of this particular isotope?

OpenStudy (anonymous):

12C: 12g 1mol actual: 3.5g xmol x=3.5/12=something (mol) number of atom = something*Avogadro number

OpenStudy (anonymous):

Well there are many ways of doing this, x=3.5/12= unknown thats stands for (mol) but really 3.5g is xmol number of it=unknown. Do you get it?

OpenStudy (anonymous):

Kind of. So for 3.5/12= 482.4?

OpenStudy (anonymous):

These are the options I have 2.11 x 1024 atoms 2.07 x 1024 atoms 8.01 x 1023 atoms 1.52 x 1023 atoms 1.76 x 1023 atoms

OpenStudy (anonymous):

Would it be the first one?

OpenStudy (anonymous):

Yes that is close enough

OpenStudy (anonymous):

This is the answer 1.76 x 1023 atoms.

OpenStudy (anonymous):

the last one is right. You calculate x then take the result times avogadro constant

OpenStudy (anonymous):

Can you guys help me out with this one? A sample of (N2H5)2C3H4O4 contains 1.084 ×1024 carbon atoms. How many moles of hydrogen atoms are in the same sample?

OpenStudy (anonymous):

Yeah, I got that.

OpenStudy (anonymous):

It's 10^24*

OpenStudy (anonymous):

I know man lol

OpenStudy (anonymous):

this is dead easy. can you count how many H and how many C in the formula?

OpenStudy (anonymous):

14 H and 3 C

OpenStudy (anonymous):

good, now do some rough work H C 14 3 x atoms 1.084 *10^24 atoms so you can find x=1.084*10^24*14/3=blabla number of mol of H = blabla/Avogadro constant

OpenStudy (anonymous):

~8.4 mol I'd say

OpenStudy (anonymous):

That's what I got but it's not in the choices I have. These are the choices! 34.7 g 1.56 × 10-21 g 937 g 6.43 × 1020 g 31.2 g

OpenStudy (anonymous):

those choices are in g, not in mol, I doubt that's not the right choices for this qs

OpenStudy (anonymous):

I think they are because it's online through a school program.

OpenStudy (anonymous):

Do I have to convert it to mol to grams now?

OpenStudy (anonymous):

Never mind, sorry. I mixed up a question. This is the right answer.

OpenStudy (anonymous):

lol so it's correct yeah? my brain is not really getting old lol

OpenStudy (anonymous):

Do you mind helping me with more?

OpenStudy (anonymous):

Haha...

OpenStudy (anonymous):

ok, just 1 more.

OpenStudy (anonymous):

The mass of 5.20 mol of glucose, C6H12O6 is? A-34.7 g B-1.56 × 10-21 g C-937 g D-6.43 × 1020 g E-31.2 g

OpenStudy (anonymous):

I get 482.2g

OpenStudy (anonymous):

It's 482.4g. But it doesn't make sense.

OpenStudy (anonymous):

you need to find the mass formula for glucose first. 6mols of C, 12mols H, 6mols O MC=12 MH=1 MO=16 so 6*12+12*1+6*16=180(g/mol) there are 5.2 mols so mass m=180*5.2=936g

OpenStudy (anonymous):

Oh... ok got it. I'm making harder than it is :P

OpenStudy (anonymous):

you know it's good to remember the mass of 1 mol of some common substance. I've done these in secondary school and I still remember them now. I'd say that's big advantage. e.g H=1, O=16, N=14, Cl=35.5, C=12... those commonly used you should remember. Save you lots of time in exam.

OpenStudy (anonymous):

I actually do remember these basic ones. Thanks for the tip. If you have any other tips, shoot'em.

OpenStudy (anonymous):

lol I wasn't that great at chemistry though. remember this too K - Na - Ca - Mg -Al - Zn - Fe- Ni - Sn - Pb - H - Cu - Hg - Ag- Pt - Au That's extremely helpful when you deal with acid and base And always try to make a table of information first, make the problem clearer. And don't get scared of those Chem teachers. When they die, they'll just turn into HCON, no more no less =))

OpenStudy (anonymous):

good luck. I'm off to bed.

OpenStudy (anonymous):

Thanks a lot :)

OpenStudy (anonymous):

no problem man